# LinkNER: Linking Local Named Entity Recognition Models to Large Language Models using Uncertainty

### v2

@@ Abstract.
Named Entity Recognition (NER) serves as a fundamental task in natural language understanding, bearing direct implications for web content analysis, search engines, and information retrieval systems. Fine-tuned NER models exhibit satisfactory performance on standard NER benchmarks. However, due to limited fine-tuning data and lack of knowledge, it performs poorly on unseen entity recognition. As a result, the usability and reliability of NER models in web-related applications are compromised. Instead, Large Language Models (LLMs) like GPT-4 possess extensive external knowledge, but research indicates that they lack specialty for NER tasks. Furthermore, non-public and large-scale weights make tuning LLMs difficult. To address these challenges, we propose a framework that combines small fine-tuned models with LLMs (LinkNER) and an uncertainty-based linking strategy called RDC that enables fine-tuned models to complement black-box LLMs, achieving better performance. We experiment with both standard NER test sets and noisy social media datasets. LinkNER enhances NER task performance, notably surpassing SOTA models in robustness tests. We also quantitatively analyze the influence of key components like uncertainty estimation methods, LLMs, and in-context learning on diverse NER tasks, offering specific web-related recommendations.
@@ 4.1. LinkNER Workflow
The entity recognition of each text contains three steps, and the workflow of LinkNER is shown in Figure 3. At step 1⃝, the test sentence is fed into the local model, which outputs the entity recognition result and the corresponding uncertainty probability, denoted as $y=[(\mathrm{x}^{e},\mathrm{x}^{d},\mathrm{label}),u]$. At step 2⃝, the uncertainty of the local model output is compared with a predefined threshold $\tau$. Entities that are greater than the threshold $\tau$ are detected as uncertain, and sent to the LLM for further classification. Entities that are less than the threshold $\tau$ do not need to undergo refinement. Finally, at step 3⃝, the outputs of the local model and LLM are integrated into the final result. Then, we explain in detail how the local model performs entity recognition and detection, and how the LLM refines the results from the local model.
@@ Appendix C Implementation Details
Dataset
Label description
CoNLL’03 CoNLL’03-Typos CoNLL’03-OOV WikiGold
”Here is the entity class information: Person: This category includes names of persons, such as individual people or groups of people with personal names. Organization: The organization category consists of names of companies, institutions, or any other group or entity formed for a specific purpose. Location: The location category represents names of geographical places or landmarks, such as cities, countries, rivers, or mountains. Miscellaneous: The miscellaneous category encompasses entities that do not fall into the above three categories. This includes adjectives, like Italian, and events, like 1000 Lakes Rally, making it a very diverse category.”
@@
”Here is the entity class information: PERSON: People, including fictional; ORGANIZATION: Companies, agencies, institutions, etc. GPE: Countries, cities, states; DATE: Absolute or relative dates or periods; NORP: Nationalities or religious or political groups; CARDINAL: Numerals that do not fall under another type; TIME: Times smaller than a day; LOC: Non-GPE locations, mountain ranges, bodies of water; FACILITY:Buildings, airports, highways, bridges, etc; PRODUCT: Vehicles, weapons, foods, etc. (Not services); WORK_OF_ART: Titles of books, songs, etc; MONEY: Monetary values, including unit; ORDINAL: “first”, “second”, etc; QUANTITY: Measurements, as of weight or distance, etc; EVENT: Named hurricanes, battles, wars, sports events, etc; PERCENT: Percentage (including “%”), etc; LAW: Named documents made into laws, etc; LANGUAGE: Any named language, etc.”
@@
”Here is all entity class information: Person: Names of people (e.g. Virginia Wade). Don’t mark people that don’t have their own name. Include punctuation in the middle of names. Fictional people can be included, as long as they’re referred to by name (e.g. Harry Potter). Location: Names that are locations (e.g. France). Don’t mark locations that don’t have their own name. Include punctuation in the middle of names. Fictional locations can be included, as long as they’re referred to by name (e.g. Hogwarts). Corporation: Names of corporations (e.g. Google). Don’t mark locations that don’t have their own name. Include punctuation in the middle of names. Product: Name of products (e.g. iPhone). Don’t mark products that don’t have their own name. Include punctuation in the middle of names. Fictional products can be included, as long as they’re referred to by name (e.g. Everlasting Gobstopper). It’s got to be something you can touch, and it’s got to be the official name. Creative-work: Names of creative works (e.g. Bohemian Rhapsody). Include punctuation in the middle of names. The work should be created by a human, and referred to by its specific name. Group: Names of groups (e.g. Nirvana, San Diego Padres). Don’t mark groups that don’t have a specific, unique name, or companies (which should be marked corporation).”
@@
”Here is the entity class information: For polysemous entities, our guidelines instructed annotators to assign the entity class that corresponds to the correct entity class in the given context. For example, in “We’re driving to Manchester”, Manchester is a Location, but in “Manchester are in the final tonight”, it is a sports club – an Organization. Special attention is given to username mentions. Where other corpora have blocked these out or classified them universally as Person, our approach is to treat these as named entities of any potential class. For example, the account belonging to the M̈anchester United football clubẅould be labeled as an Organization. Other: The Other category encompasses entities that do not fall into the above three categories. This includes adjectives, like Italian, and events, like 1000 Lakes Rally, making it a very diverse category.”
@@
”Here is the entity class information: The composition of the Protein category includes protein molecules, complexes, etc. The DNA is Deoxyribonucleic Acid. The RNA is Ribonucleic Acid. The Cell_type category refers to the Cell type in nature. The Cell_line category refers to the Cell line in artificial.”

### v3

@@ Abstract.
Named Entity Recognition (NER) serves as a fundamental task in natural language understanding, bearing direct implications for web content analysis, search engines, and information retrieval systems. Fine-tuned NER models exhibit satisfactory performance on standard NER benchmarks. However, due to limited fine-tuning data and lack of knowledge, it performs poorly on unseen entity recognition. As a result, the usability and reliability of NER models in web-related applications are compromised. Instead, Large Language Models (LLMs) like GPT-4 possess extensive external knowledge, but research indicates that they lack specialty for NER tasks. Furthermore, non-public and large-scale weights make tuning LLMs difficult. To address these challenges, we propose a framework that combines small fine-tuned models with LLMs (LinkNER) and an uncertainty-based linking strategy called RDC that enables fine-tuned models to complement black-box LLMs, achieving better performance. We experiment with both standard NER test sets and noisy social media datasets. LinkNER enhances NER task performance, notably surpassing SOTA models in robustness tests. We also quantitatively analyze the influence of key components like uncertainty estimation methods, LLMs, and in-context learning on diverse NER tasks, offering specific web-related recommendations. Code is available at https://github.com/LinkNER
@@ 4.1. LinkNER Workflow
[…]
@@ Appendix C Implementation Details
Dataset Label description
CoNLL’03 CoNLL’03-Typos CoNLL’03-OOV WikiGold ”Here is the entity class information: Person: This category includes names of persons, such as individual people or groups of people with personal names. Organization: The organization category consists of names of companies, institutions, or any other group or entity formed for a specific purpose. Location: The location category represents names of geographical places or landmarks, such as cities, countries, rivers, or mountains. Miscellaneous: The miscellaneous category encompasses entities that do not fall into the above three categories. This includes adjectives, like Italian, and events, like 1000 Lakes Rally, making it a very diverse category.”